OpenStudy (anonymous):

A bag contains 45 marbles of various colors. Jon draws 60 times, each time replacing the marble before drawing again. He gets a total of 12 blue marbles. Which is best prediction of how many blue marbles are in the bag? A. 5 B. 9 C. 12 D. 15

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

@it_meh_miranda

OpenStudy (anonymous):

hey sry i didnt help before my os wasnt working

OpenStudy (hugsnotughs):

Oh gosh I hated these problems in middle school.

OpenStudy (kropot72):

The experimental probability of drawing a blue marble is 12/60 = 1/5. There are 45 marbles in the bag, so the expected number of blue marbles is: \[45\times\frac{1}{5}=you\ can\ calculate\]